# Quantum Mixing and Benjamini–Schramm Convergence of Hyperbolic Surfaces

### v1

@@ Abstract.
We study compact hyperbolic surfaces and multiplicative observables, establishing a large-scale analogue of Zelditch’s quantum mixing theorem with hypotheses that hold for both arithmetic and Weil-Petersson random surfaces of large genus. This complements the large-scale quantum ergodicity theorems of Le Masson and Sahlsten, which themselves serve as large-scale analogues of the quantum ergodicity theorem of Shnirelman, Zelditch, and de Vèrdiere, thereby providing a more complete picture of the asymptotic behaviour of observables in the large-scale limit. Our approach does not rely on ball averaging operators or Nevo’s ergodic theorem. Instead, we introduce a new method based on the hyperbolic wave equation and the quantitative exponential mixing of the geodesic flow established by Ratner and Matheus.
The research conducted here is supported by the Vilho, Yrjö and Kalle Väisälä Foundation, Research council of Finland project: Quantum chaos on large and many body systems grant numbers 347365, 353738, and Aalto University
@@ 1.1. Background
In quantum systems, the state of the system can be described in terms of the eigenfunctions of its energy operator. Their evolution is governed by the Schrödinger equation, the quantum analogue of Newton’s equations of motion. Observables correspond to measurable quantities of the system. Understanding the interplay between the energy operator, its eigenfunctions, and the observables leads to a deeper understanding of the dynamics of the system. A central topic in quantum chaos is the study of localization and delocalization phenomena for the eigenfunctions of the energy operator. In this paper, the Hamiltonian system is assumed to be the geodesic flow $\phi_{t}$ of a compact connected Riemannian manifold $(M,g)$ and the quantum Hamiltonian is taken to be $\sqrt{-\Delta}$, where $\Delta$ is the Laplace-Beltrami operator on the manifold. For more background and a better overview, we refer to the following surveys: [Mar06, Zel05, Sar11, Zel19, Ana22].
Consider a compact Riemannian manifold $(M,g)$ and the quantum Hamiltonian $\sqrt{-\Delta}$ with eigenvalues $\{\mu_{j}=\sqrt{\lambda_{j}}\}$ and eigenfunctions $\{\psi_{j}\}$, where $\{\lambda_{j}\}$ and $\{\psi_{j}\}$ are the eigenvalues and eigenfunctions of the Laplace-Beltrami operator $\Delta$. Let $A$ be a zeroth-order pseudodifferential operator representing an observable. We are interested in the diagonal elements $\langle\psi_{j},A\psi_{j}\rangle$, which correspond to the expected value of $A$ in the state $\psi_{j}$, and in the off-diagonal elements $\langle\psi_{j},A\psi_{k}\rangle$, which represent transition amplitudes from $\psi_{j}$ to $\psi_{k}$ induced by the observation of $A$.
Quantum ergodicity theorem (Shnirelman [Shn74], Zelditch [Zel87], and Colin de Verdière [Col85]) is a cornerstone result of quantum chaos and serves as a quantum analogue of the classical ergodic theorem. For an ergodic compact Riemannian manifold, it states that almost all subsequences of eigenfunctions corresponding to increasing eigenvalues satisfy
@@
where $\omega(A)$ denotes the average of the principal symbol $\sigma_{A}$ of $A$ (the classical counterpart of observable $A$) over the unit cotangent bundle of $M$. Roughly speaking, the theorem asserts that the probability distributions induced by the eigenfunctions converge to the uniform distribution in the weak sense of testing the distributions against observables.
Shnirelman’s theorem concerns the diagonal elements of observables. Zelditch [Zel90, Zel96b] extended it to the off-diagonal elements, given that the compact Riemannian surface is now weak mixing, proving that for any $\epsilon>0$, there exists $\delta(\epsilon)>0$ such that for all $\tau\in\mathbb{R}$,
@@
Informally, this means that the summed transition amplitudes between states separated approximately by energy $\tau$ decay to zero for almost all subsequences of eigenfunctions as the eigenvalue increases.
A unified formulation encompassing both of these results can be given following the works of multiple authors: [Shn74, Col85, Zel87, Laz93, Zel90, GL93, Zel96a, Zel96b, ZZ96, Sun97]. The version below adapts Theorems 4.1 and 4.4 from [Zel05].
@@
Let $(M,g)$ be a compact Riemannian manifold (possibly with a boundary) with quantum Hamiltonian $\sqrt{-\Delta}$ with eigenvalues $\{\mu_{j}=\sqrt{\lambda_{j}}\}$ and eigenvalues $\{\psi_{j}\}$, where $\{\lambda_{j}\}$ and $\{\psi_{j}\}$ are the eigenvalues and eigenfunctions of the Laplace-Beltrami operator $\Delta$. Let $A$ be a zeroth-order pseudodifferential operator. Consider the following properties:
@@
    For every $\epsilon>0$ there exists $\delta(\epsilon)>0$ such that
    $\limsup_{\mu\to\infty}\ \frac{1}{\#\{j\ :\ \mu_{j}<\mu\}}\sum_{\lambda_{j}<\lambda}\sum_{\begin{subarray}{c}k:\mu_{k}<\mu,\\ k\neq j,\\ |\mu_{j}-\mu_{k}|<\delta(\epsilon)\end{subarray}}\left|\langle\psi_{j},A\psi_{k}\rangle\right|^{2}<\epsilon,$
@@
Then the geodesic flow $\varphi_{t}$ is ergodic on $(S^{*}M,d\mu)$ if and only if properties (i) and (ii) hold. It is weakly mixing on $(S^{*}M,d\mu)$ if and only if properties (i), (ii), and (iii) hold. Here $S^{*}M$ denotes the unit tangent bundle of $M$ and $d\mu$ is the Liouville measure on $S^{*}M$.
@@
Our focus will be on the compact connected hyperbolic surfaces, which are inherently chaotic due to their (constant) negative curvature. In this context, the following theorem of Ratner [Rat87], with quantitative bounds due to Matheus [Mat13], will be central. The theorem is stated here in a slightly modified form using only the physical space observables.
@@
Let $X$ be a compact connected hyperbolic surface with geodesic flow $\phi_{t}$ and with $\lambda_{1}(X)$ being the smallest eigenvalue of the Laplace-Beltrami operator $\Delta$. Let $f,g\in L^{2}(X)$. Then for all $t\geq 0$,
$\bigg|\big\langle f\circ\varphi_{t},g\big\rangle_{L^{2}(T^{1}X)}-\int_{T^{1}X}f\,dx\int_{T^{1}X}g\,dx\bigg|\leq 11\,e^{\beta(\lambda_{1}(X))}\,(1+t)\,e^{-\beta(\lambda_{1}(X))t}\,\|f\|_{2}\,\|g\|_{2},$
@@
and where $f(x)=f(x,\theta)$ and $g(x)=g(x,\theta)$ are viewed as functions on $T^{1}X$, the unit tangent bundle.
@@
The more general phase space version (see [Rat87, Mat13]) implies that the geodesic flow on a compact connected hyperbolic surface is weak mixing with exponential decay which further implies ergodicity. By Theorem 1.1, a compact connected hyperbolic surface satisfies properties (i)(iii), and hence is both quantum ergodic and quantum weak mixing.
Remark 1.3.
From Theorem 1.1, we see that neither the $2$-torus nor the $2$-sphere is quantum ergodic or quantum weakly mixing. However, if we restrict the observables to be taken from the collection of the physical-space observables, i.e. the multiplicative operators corresponding to functions, then a $2$-torus with irrational aspect ratio is quantum ergodic (though still not quantum weak mixing). The $2$-sphere remains neither quantum ergodic nor quantum weak mixing. These examples can be analyzed explicitly due to the availability of closed-form expressions for their eigenfunctions and eigenvalues. An illustarting example showing this for a discrete circle with $N$ elements can be acquired using observable $a(x)=e^{\frac{2\pi i\,x}{N}}$; this was pointed out by Mostafa Sabri. This idea should translate to the $2$-dimensional surfaces as well. We see that the choice of observables matters.
@@
These phenomena on Riemannian surfaces naturally motivate the search for analogous results in other contexts, such as discrete settings. A pseudo-analogue of the above corollary to Theorem 1.1 was obtained for certain finite regular graphs by Anantharaman and Le Masson [AL15]. Their result, however, is not a strict analogue, since the high-energy limit is not meaningful for finite graphs, which possess only finitely many eigenvalues. Instead, they considered a sequence $(G_{n})$ of graphs with increasing size $n$ and a fixed energy window $I$, as opposed to a fixed structure with increasing energy as in Shnirelman’s theorem. A special case of their result states that, under suitable assumptions, if $(a_{n})$ is a sequence of diagonal observables acting on functions on the vertices and if $(\Delta_{n})$ is a sequence of graph Laplacians each with eigenvalues $\{\lambda_{j}^{(n)}\}$ and eigenfunctions $\{\psi_{j}^{(n)}\}$, then
@@
$\frac{1}{\#\{j:\lambda_{j}^{(n)}\in I\}}\sum_{\lambda_{j}^{(n)}\in I}\left|\left\langle\psi_{j}^{(n)},a_{n}\psi_{j}^{(n)}\right\rangle-\frac{\operatorname{Tr}(a_{n})}{n}\right|^{2}\xrightarrow{n\to\infty}0.$
@@
As in Shnirelman’s theorem, the averages of the expectation values converge to the uniform average. We refer to such results, in which the structure grows while energy remains bounded, as large-scale results, in contrast to the original results in which the energy frows. We refer to the latter results as large-energy results. The large-scale quantum ergodicity of graphs and related theorems is an active field with a vigorous interplay with the continuous settings. For more realted results in the discrete graph setting, see [DP12, BL13, EKYY13, TVW13, AL15, Gei15, BLL16, Ana17, AS17, BHKY17, BKY17, AS19, BS19, BHY19, McK22, MS23]. An especially impressive result establsihed in [BHKY17, BHY19, BKY17] shows a large-scale analogue of the quantum unique ergodicity conjecture of Rudnick and Sarnak [RS94] for random regular graphs; the randomization effectively suppresses pathological cases helping to establish stronger results compared to the deterministic setting.
@@
Since the large-scale graph quantum ergodicity theorem is not a true analogue of Shnirelman’s theorem, it is natural to seek a corresponding surface result highlighting the distinctions from the original large-energy setting. Such an analogue for hyperbolic surfaces was established by Le Masson and Sahlsten [LS17] (see Subsection 1.2 for details). Their result resembles closely that of Anantharaman and Le Masson, with minor modifications reflecting the transition from discrete graphs to continuous surfaces.
Despite the distinct nature of the two limits, there exists a heuristic correspondence between the large-energy and large-scale settings. In the large-energy regime, as the energy grows, the eigenfunctions exhibit shorter wavelengths, resolving increasingly fine structures. In the large-scale regime, as the structure grows, finer details become accessible within a fixed energy window. Hence, these two asymptotic regimes may be regarded as analogous in spirit.
@@
To our knowledge, the only existing large-scale off-diagonal results establishing analogues of properties (i), (ii), and (iii) of Theorem 1.1 (see Subsection 1.2) are those of Cipolloni, Erdős, and Schröder [CES21] for Wigner matrices. Their setting resembles that of Anantharaman and Le Masson, but with graphs replaced by random matrices and observables by general matrices. Their result is remarkably strong as they require no averaging: they prove that
@@
$\left|\left\langle\psi_{j}^{(n)},A_{n}\psi_{k}^{(n)}\right\rangle-\delta_{jk}\tfrac{\operatorname{Tr}(A_{n})}{n}\right|$
@@
decays to zero at the optimal and uniform rate of $n^{-1/2}$ for all $j,k$, thereby establishing a large-scale analogue of the quantum unique ergodicity conjecture of Rudnick and Sarnak [RS94], and in fact extending it to the off-diagonal elements. This result is relevant as the random matrix theory is closely related to the quantum ergodicity on graphs and surfaces, especially if Berry’s conjecture (see [Ber77]) holds. For more related results for random matrices, see [ESY09a, ESY09b, BY17].
@@
A new upcoming result by Bordenave, Letroit, and Sabri [BLS] establishes a quantum ergodicy and quantum mixing results on growing Schreier graphs where even the quantum ergodicy was previously unknown. To our knowledge, this is first mixing result in the graph setting. This is especially interesting as graphs are intuitively closer to the original setting on manifolds when compared to random matrices.

### v2

@@ Abstract.
We study compact hyperbolic surfaces and multiplication observables, establishing a large-scale analogue of Zelditch’s quantum mixing theorem with hypotheses that hold for both arithmetic and Weil–Petersson random surfaces of large genus. This complements the large-scale quantum ergodicity theorems of Le Masson and Sahlsten, which themselves are large-scale analogues of the quantum ergodicity theorem of Shnirelman, Zelditch, and Colin de Verdière, thereby providing a more complete picture of the asymptotic behavior of observables in the large-scale limit. Our approach does not rely on the ball averaging operator or Nevo’s ergodic theorem. Instead, we introduce a new method based on the hyperbolic wave equation and the quantitative exponential mixing of the geodesic flow established by Ratner and Matheus.
The research conducted here is supported by the Vilho, Yrjö and Kalle Väisälä Foundation, Research Council of Finland project: Quantum chaos on large and many body systems, grant numbers 347365, 353738, and Aalto University.
@@ 1.1. Background
In quantum systems, the state of the system can be described in terms of the eigenfunctions of its energy operator. Their evolution is governed by the Schrödinger equation, the quantum analogue of Newton’s equations of motion. Observables correspond to measurable quantities of the system. Understanding the interplay between the energy operator, its eigendata (eigenvalues and eigenfunctions), and the observables leads to a deeper understanding of the dynamics of the system. A central topic in quantum chaos is the study of localization and delocalization phenomena of the eigenfunctions of the energy operator. Here, the Hamiltonian system is assumed to be the geodesic flow $\varphi_{t}$ of a compact connected Riemannian manifold $(M,g)$ and the quantum Hamiltonian is taken to be $\sqrt{-\Delta}$, where $\Delta$ is the Laplace–Beltrami operator of the manifold. For more background and a better overview, we refer to the following surveys: [ZEL05, SAR11, ZEL19, ANA02, MAR02].
Consider a compact Riemannian manifold $(M,g)$ and the quantum Hamiltonian $\sqrt{-\Delta}$ with eigenvalues $\{\mu_{j}=\sqrt{\lambda_{j}}\}$ and eigenfunctions $\{\psi_{j}\}$, where $\{\lambda_{j}\}$ and $\{\psi_{j}\}$ are the eigenvalues and eigenfunctions of the Laplace–Beltrami operator $\Delta$. Let $A$ be a zeroth-order pseudo-differential operator representing an observable. We are interested in the diagonal average $\langle\psi_{j},A\psi_{j}\rangle$, which corresponds to the expectation value of $A$ in the state $\psi_{j}$, and in the off-diagonal average $\langle\psi_{j},A\psi_{k}\rangle$, which represents the transition amplitude from $\psi_{j}$ to $\psi_{k}$.
Quantum ergodicity theorem (Shnirelman [ŠNI74], Zelditch [ZEL87], and Colin de Verdière [COL85]) is a cornerstone result of quantum chaos and serves as a quantum analogue of the classical ergodic theorem. For an ergodic compact Riemannian manifold, it states that almost all subsequences of eigenfunctions corresponding to increasing eigenvalues satisfy that
@@
where $\omega(A)$ denotes the average of the principal symbol $\sigma_{A}$ of $A$ (the classical counterpart of observable $A$) over the unit co-tangent bundle of $M$. Roughly speaking, the theorem asserts that the probability distributions induced by the eigenfunctions converge to the uniform distribution in the weak sense of testing the distributions against observables.
The quantum ergodicity theorem focuses on the diagonal averages of the observables. Zelditch [ZEL90, ZEL96b] extended it to the off-diagonal averages, given that the geodesic flow on the compact Riemannian surface is now weak mixing, proving that for any $\epsilon>0$, there exists $\delta(\epsilon)>0$ such that for all $\tau\in\mathbb{R}$
@@
This means that almost all subsequences of eigenfunctions with increasing eigenvalues, $(\psi_{j_{l}})$, satisfy that the transition amplitudes into a small energy window of separation $\tau$ from the eigenvalue of $\psi_{j_{l}}$ can be made arbitrarily small as $l\to\infty$ by choosing a small enough energy window.
A unified formulation encompassing both of these results can be given following the works of multiple authors: [ŠNI74, COL85, ZEL87, ZEL90, GL93, LAZ93, ZEL96a, ZEL96b, ZZ96, SUN97]. The version below adapts Theorems 4.1 and 4.4 from [ZEL05].
@@
Let $(M,g)$ be a compact Riemannian manifold (possibly with a boundary) with quantum Hamiltonian $\sqrt{-\Delta}$ together with eigenvalues $\{\mu_{j}=~\sqrt{\lambda_{j}}\}$ and eigenfunctions $\{\psi_{j}\}$, where $\{\lambda_{j}\}$ and $\{\psi_{j}\}$ are the eigenvalues and eigenfunctions of the Laplace–Beltrami operator $\Delta$. Let $A$ be a zeroth-order pseudo-differential operator. Consider the following properties:
@@
    For every $\epsilon>0$, there exists $\delta(\epsilon)>0$ such that
    $\limsup_{\mu\to\infty}\ \frac{1}{\#\{j:\mu_{j}<\mu\}}\sum_{j:\mu_{j}<\mu}\sum_{\begin{subarray}{c}k:\mu_{k}<\mu,\\ k\neq j,\\ |\mu_{j}-\mu_{k}|<\delta(\epsilon)\end{subarray}}\left|\langle\psi_{j},A\psi_{k}\rangle\right|^{2}<\epsilon,$
@@
    For every $\epsilon>0$, there exists $\delta(\epsilon)>0$ such that for all $\tau\in\mathbb{R}$,
@@
Then the geodesic flow $\varphi_{t}$ is ergodic on $(S^{*}M,d\mu)$ if and only if Properties (I) and (II) hold. It is weakly mixing on $(S^{*}M,d\mu)$ if and only if Properties (I)–(III) hold. Here $S^{*}M$ denotes the unit co-tangent bundle of $M$ and $d\mu$ is the Liouville measure on $S^{*}M$.
@@
Our focus will be on compact connected hyperbolic surfaces, which are inherently chaotic due to their (constant) negative curvature. In this context, the following theorem of Ratner [RAT87], with quantitative bounds due to Matheus [MAT13], will be central. The theorem is stated here in a slightly modified form using only the physical space observables.
@@
Let $X$ be a compact connected hyperbolic surface with geodesic flow $\varphi_{t}$ and with $\lambda_{1}(X)$ being the second smallest eigenvalue of the Laplace–Beltrami operator $\Delta$. Let $f,g\in L^{2}(X)$. Then for all $t\geq 0$
$\bigg|\big\langle f\circ\varphi_{t},g\big\rangle_{L^{2}(S^{*}X)}-\int_{S^{*}X}f\,dx\int_{S^{*}X}g\,dx\bigg|\leq 11\,e^{\beta(\lambda_{1}(X))}\,(1+t)\,e^{-\beta(\lambda_{1}(X))t}\,\|f\|_{2}\,\|g\|_{2},$
@@
and where $f(x)=f(x,\theta)$ and $g(x)=g(x,\theta)$ are viewed as functions on $S^{*}X$, the unit co-tangent bundle.
@@
The more general phase space version (see [RAT87, MAT13]) implies that the geodesic flow on a compact connected hyperbolic surface is weak mixing, which further implies ergodicity. By Theorem 1.1, a compact connected hyperbolic surface satisfies Properties (I)(III), and hence is both quantum ergodic and quantum mixing.
From Theorem 1.1, it follows that neither the $2$-torus nor the $2$-sphere is quantum ergodic or quantum mixing, since their geodesic flows are neither ergodic nor weakly mixing. However, if one restricts the class of observables to physical-space observables, that is, multiplication operators corresponding to functions, then all flat $2$-tori satisfy Property (I); see [MR12] for details. This illustrates that the choice of observables plays a crucial role. On the other hand, it is an interesting question whether Properties (II) and (III) hold for flat $2$-tori. For further discussion, see Appendix A.
@@
These phenomena on Riemannian surfaces naturally motivate the search for analogous results in other settings, for example, the discrete setting. Anantharaman and Le Masson [AL15] established that, for certain finite graphs, the analogue of Property (I) holds. Their result, however, is not a strict analogue, since finite graphs have only finitely many eigenvalues. Instead, they considered a sequence $(G_{n})$ of graphs with increasing size $n$ and a fixed energy window $I$; as the graph grows, the number of eigenvalues in the energy window increases. This differs from the quantum ergodicity theorem, in which one considers a fixed structure at increasing energy levels, so that all eigenvalues below a given energy and the corresponding eigenfunctions are considered.
A special case of the theorem of Anantharaman and Le Masson [AL15] states that if $(a_{n})$ is a uniformly bounded sequence of diagonal matrices acting on functions on the vertices of the graph, and if $(\Delta_{n})$ is a sequence of graph Laplacians, each with eigenvalues $\{\lambda_{j}^{(n)}\}$ and eigenfunctions $\{\psi_{j}^{(n)}\}$, then under certain conditions (analogous to Properties (BSC) and (EXP) introduced in Subsection 1.2) on the structure of the graphs, we have
@@
$\lim_{n\to\infty}\frac{1}{\#\{j:\lambda_{j}^{(n)}\in I\}}\sum_{j:\lambda_{j}^{(n)}\in I}\left|\left\langle\psi_{j}^{(n)},a_{n}\psi_{j}^{(n)}\right\rangle-\frac{\operatorname{Tr}(a_{n})}{n}\right|^{2}=0.$
As in the quantum ergodicity theorem, the averages of the expectation values converge to the average of the observable with respect to the uniform distribution.
@@
We refer to such results in which the structure grows while the energy remains bounded as large-scale results, in contrast to the original results in which the energy grows. We refer to the latter results as large-energy results. The study of the large-scale quantum ergodicity of graphs and related problems is an active field with a vigorous interplay with the continuous setting. For more related results in the discrete graph setting, see [DP12, BL13, EKY+13, TVW13, AL15, GEI15, BLL16, ANA17, AS17, BHK+17a, BKY17b, AS19, BS19, BHY19, MCK22, MS23]. An impressive result established in [BHK+17a, BHY19, BKY17b] shows an average-free large-scale analogue of the large-scale quantum ergodicity theorem of Anantharaman and Le Masson [AL15] for random regular graphs. This resembles the quantum unique ergodicity conjecture of Rudnick and Sarnak [RS94] in the random setting. However, the observable in [BHK+17a, BHY19, BKY17b] must be chosen independently of the graph. Nevertheless, the randomization effectively suppresses pathological cases, helping to establish stronger results when compared to the deterministic setting.
@@
Since the large-scale quantum ergodicity theorem for graphs by Anantharaman and Le Masson [AL15] differs from the large-energy quantum ergodicity theorem of Shnirelman [ŠNI74], Zelditch [ZEL87], and Colin de Verdière [COL85], it is natural to seek a corresponding surface result for the large-scale graph theorem that would highlight distinctions when compared to the original large-energy quantum ergodicity theorem. Such an analogue for hyperbolic surfaces was established by Le Masson and Sahlsten [LS17] (see Subsection 1.2 for more details). Their result closely resembles that of Anantharaman and Le Masson in that they prove a similar analogue of Property (I) of the original quantum ergodicity theorem; minor differences reflect the transition from discrete graphs to continuous surfaces.
It is tempting to try to equate the large-energy setting with the large-scale setting. Heuristically, one might say that in the large-energy regime we use higher and higher energies to resolve finer details, whereas in the large-scale setting we zoom into the structure to capture finer features. However, the large-scale setting appears to be much more sensitive to the underlying structure of the manifold or graph. The large-energy quantum ergodicity theorem says that all Riemannian manifolds for which the geodesic flow is ergodic are quantum ergodic and that those for which the geodesic flow is weak mixing are quantum mixing; this is open for the variable-curvature large-scale setting. Moreover, in the case of graphs, the semiclassical measure is not the uniform measure, so it is not expected to be uniform either on manifolds of variable negative curvature, in contrast to the large-energy regime, which does not see this subtlety.
@@
To our knowledge, the only previously existing large-scale off-diagonal results establishing analogues of Properties (I)(III) of Theorem 1.1 (see Subsection 1.2 for the properties) are those of Cipolloni, Erdős, and Schröder [CES21] for Wigner matrices. Their setting resembles that of Anantharaman and Le Masson [AL15], where graphs are replaced by random matrices. Their result is remarkably strong, as they require no averaging: they prove that
@@
$\lim_{n\to\infty}\left|\left\langle\psi_{j}^{(n)},A_{n}\psi_{k}^{(n)}\right\rangle-\delta_{jk}\tfrac{\operatorname{Tr}(A_{n})}{n}\right|=0;$
@@
the decay to 0 has the optimal and uniform rate of $n^{-1/2}$ for all $j,k$, where $n\times n$ matrices are considered with $n$ growing. Thereby Cipolloni, Erdős, and Schröder [CES21] establish a result that resembles the quantum unique ergodicity conjecture of Rudnick and Sarnak [RS94] in the large-scale setting; moreover, they extend it to the off-diagonal case as well. We remark that, as in [BHK+17a, BHY19, BKY17b], the observable is chosen independently of the matrix. This result is relevant, as random matrix theory is closely related to the study of quantum ergodicity and related problems on graphs and surfaces, especially if Berry’s conjecture holds (see [BER77]). For more related results for random matrices, see [ESY09a, ESY09b, BY17].
@@
Studying the off-diagonal setting is timely: a new result by Bordenave, Letrouit, and Sabri [BLS26] establishes quantum ergodicity and quantum mixing results on growing Schreier graphs, where even quantum ergodicity was previously unknown. To our knowledge, this is the first mixing result in the graph setting. This is especially interesting, as the graphs are intuitively closer to the original setting on manifolds when compared to random matrices.